Grade 12th passInorganic Chemistry

If the weight of metal chloride is x gram containing y gram of metal, the equivalent weight of metal will be?

Ruchika Sisodiya
9 Years agoGrade 12th pass

2 Answers

aneesu
9 Years ago
answer for this question is y(35.5)/{x-y]. this can be answered by equating equivalent of metal and metal cloride .hope this will do
 
Pooja Kumari
9 Years ago
Let Metal Chloride be MCla(×-y)/35.5a = x/(x/a)Equivalent weight of metal is y/a = y÷((x-y)/35.5)^(1/2)Answer: y*{(35.5/(x-y))^(1/2)}